OpenStudy (anonymous):

What is the energy of a photon of light having a wavelength of 4.5 x 10-7 meters?

OpenStudy (cuanchi):

You have to use the equation for energy E=hc/λ h = 6.626 × 10^( -34) joule·s (Plank's constant) c = 2.998 × 10^8 m/s (speed of the light in the vacuum) λ = wavelength (m) the answer will be in Joules, if you need in electron volts you have to use the conversion factor 1 eV = 1.602 × 10-19 J. http://www.pveducation.org/pvcdrom/properties-of-sunlight/energy-of-photon